OpenStudy (anonymous):

under which condition does sin squared x + sin squared y = 1? sorry im not sure how to type exponents on my computer

OpenStudy (sirm3d):

\[\sin^2 x + \sin^2 y=1\] compare this with a known identity \[\sin^2 x + \cos^2 x =1\] the logical solution is \[\sin^2 y=\cos^2 x\] also, \[\sin y = \cos(\pi/2 - y)=\cos(y-\pi/2)\] therefore \[\cos^2(\pi/2-y)=\cos^2 x=\cos^2(y-\pi/2)\] one solution is given by \(\pi/2-y=x\) and another solution is \(y-\pi/2=x\) there are solutions other than those i gave.